# ACING: Actor-Critic for Instruction Learning in Black-Box Large Language Models

### v1

@@ Appendix C ACING with Budget Splitting
Spelling Letters_list 0.59 (0.02) 1.00 (0.00) 0.99 (0.01) 1.00 (0.00) 1.00 (0.00)
Second_word_letter 0.00 (0.00) 0.35 (0.09) 0.39 (0.28) 0.40 (0.17) 0.70 (0.15)
Morpho-Syntax Singular_to_plural 1.00 (0.00) 0.99 (0.01) 0.95 (0.03) 0.99 (0.01) 0.95 (0.03)
Active_to_passive 1.00 (0.00) 0.98 (0.01) 1.00 (0.00) 1.00 (0.00) 1.00 (0.00)
Negation 0.79 (0.00) 0.65 (0.10) 0.58 (0.22) 0.82 (0.00) 0.71 (0.06)
Lexical Semantics Antonyms 0.79 (0.02) 0.76 (0.00) 0.84 (0.01) 0.76 (0.06) 0.74 (0.01)
Synonyms 0.14 (0.01) 0.22 (0.11) 0.19 (0.08) 0.12 (0.02) 0.13 (0.02)
Word_unscrambling 0.54 (0.00) 0.59 (0.06) 0.54 (0.02) 0.59 (0.05) 0.50 (0.07)
Phonetics Rhymes 0.59 (0.01) 0.99 (0.01) 0.36 (0.04) 0.60 (0.38) 0.57 (0.31)
Numerical Sum 0.87 (0.01) 1.00 (0.00) 0.70 (0.21) 0.81 (0.30) 0.69 (0.31)
Diff 0.00 (0.00) 1.00 (0.00) 0.93 (0.09) 0.97 (0.04) 1.00 (0.00)
Knowledge Larger_animal 0.72 (0.02) 0.63 (0.07) 0.81 (0.09) 0.86 (0.06) 0.84 (0.07)
Periodic_elements 0.99 (0.01) 0.96 (0.03) 1.00 (0.00) 0.98 (0.03) 0.98 (0.00)
Cognitive Tasks Cause_and_effect 0.44 (0.09) 0.52 (0.09) 0.55 (0.11) 0.76 (0.18) 0.69 (0.15)
Common_concept 0.03 (0.02) 0.14 (0.04) 0.09 (0.04) 0.10 (0.01) 0.19 (0.05)
Object_counting 0.30 (0.02) 0.38 (0.06) 0.40 (0.12) 0.48 (0.11) 0.41 (0.03)
Odd_one_out 0.32 (0.02) 0.57 (0.02) 0.25 (0.18) 0.59 (0.05) 0.64 (0.00)
Orthography_starts_with 0.23 (0.01) 0.41 (0.09) 0.54 (0.06) 0.54 (0.15) 0.60 (0.12)
Taxonomy_animal 0.02 (0.02) 0.67 (0.14) 0.85 (0.06) 0.53 (0.34) 0.71 (0.02)
Auto_categorization 0.31 (0.01) 0.29 (0.02) 0.07 (0.07) 0.27 (0.06) 0.29 (0.04)
Word_sorting 0.58 (0.01) 0.64 (0.05) 0.23 (0.20) 0.72 (0.02) 0.70 (0.03)
CLUE Sentence_similarity 0.00 (0.00) 0.10 (0.00) 0.00 (0.00) 0.13 (0.08) 0.13 (0.07)
Sentiment 0.90 (0.00) 0.88 (0.03) 0.88 (0.02) 0.88 (0.03) 0.89 (0.01)
Translation Num_to_verbal 0.13 (0.02) 0.99 (0.01) 1.00 (0.00) 1.00 (0.00) 0.99 (0.01)
Translation_en-de 0.83 (0.01) 0.82 (0.01) 0.77 (0.02) 0.82 (0.01) 0.82 (0.01)
Translation_en-es 0.86 (0.01) 0.67 (0.24) 0.89 (0.00) 0.86 (0.02) 0.87 (0.02)
Translation_en-fr 0.88 (0.01) 0.77 (0.06) 0.85 (0.02) 0.85 (0.02) 0.83 (0.01)
Style Informal_to_formal 0.57 (0.01) 0.48 (0.02) 0.54 (0.09) 0.44 (0.05) 0.44 (0.05)
Coding Auto_debugging 0.25 (0.00) 0.25 (0.00) 0.07 (0.07) 0.29 (0.07) 0.25 (0.00)
median score 0.49 0.66 0.64 0.76 0.71
# best-performing tasks 7 8 8 15 14
Table 5: Average test performance (and standard deviations) across 3 random seeds comparing ACING versus APE, InstructZero, and INSTINCT. The bottom rows report the median score and total number of best-performing tasks for each method.
Appendix D ACING with Different Intrinsic (Action) Dimensions
@@
# best-performing tasks 8 9 10 15 14
@@
Appendix E ACING with Different Number of Exemplars
@@
# best-performing tasks 12 23
@@
Appendix F ACING with Different White-box models
@@
# best-performing tasks 14 21
@@ Appendix G Demonstrations with Human Instructions
Spelling First_word_letter cat $\rightarrow$ c Extract the first letter of the input word. 1.00 1.00
Second_word_letter cat $\rightarrow$ a Extract the second letter of the input word. 0.96 0.59
Letters_list cat $\rightarrow$ c a t Break the input word into letters, separated by spaces. 1.00 1.00
Morpho- syntax Singular_to_plural cat $\rightarrow$ cats Convert the input word to its plural form. 1.00 1.00
Active_to_passive The artist introduced the scientist. $\rightarrow$ The scientist was introduced by the artist. Write the input sentence in passive form. 1.00 1.00
Syntax Negation Time is finite $\rightarrow$ Time is not finite. Negate the input sentence. 0.81 0.82
Lexical Semantics Antonyms won $\rightarrow$ lost Write a word that means the opposite of the input word. 0.70 0.82
Synonyms alleged $\rightarrow$ supposed Write a word with a similar meaning to the input word. 0.14 0.13
Phonetics Rhymes sing $\rightarrow$ ring Write a word that rhymes with the input word. 0.61 1.00
Knowledge Larger_animal koala, snail $\rightarrow$ koala Write the larger of the two given animals. 0.94 0.94
Semantics Cause_and_effect Sentence 1: The soda went flat. Sentence 2: The bottle was left open. $\rightarrow$ The bottle was left open. Find which of the two given cause and effect sentences is the cause. 0.97 0.92
Common_concept guitars, pendulums, neutrinos $\rightarrow$ involve oscillations. Find a common characteristic for the given objects. 0.11 0.11
Style Informal_to_formal Please call once you get there $\rightarrow$ Please call upon your arrival. Rephrase the sentence in formal language. 0.63 0.5
Numerical Sum 22 10 $\rightarrow$ 32 Sum the two given numbers. 1.00 1.00
Diff 32 22 $\rightarrow$ 10 Subtract the second number from the first. 1.00 1.00
Num_to_Verbal 26 $\rightarrow$ twenty-six Write the number in English words. 1.00 1.00
Multi- lingual Translation_en-de game $\rightarrow$ Spiel Translate the word into German. 0.81 0.84
Translation_en-es game $\rightarrow$ juego Translate the word into Spanish. 0.89 0.88
Translation_en-fr game $\rightarrow$ jeu Translate the word into French. 0.86 0.87
GLUE Sentiment The film is small in scope, yet perfectly formed. $\rightarrow$ positive Determine whether a movie review is positive or negative. 0.89 0.90
Sentence_similarity Sentence 1: A man is smoking. Sentence 2: A man is skating. $\rightarrow$ 0 - definitely not Rate the semantic similarity of two input sentences on a scale of 0 - definitely not to 5 - perfectly. 0.00 0.21
median score 0.89 0.90
average score 0.78 0.79
@@
Table 9: Classified tasks into categories from the instruction-induction datasets. For each task, we provide a corresponding demonstration, with $\rightarrow$ separating the input from the output, along with its respective human instruction as proposed in Honovich et al. (2023). We tested these instructions, report their test scores, and compare them to our best test scores using ACING with Vicuna-13B as the white-box LLM.
Appendix H Our Best Learned Instructions
@@
second_word_letter Output the first letter of each word that is a vowel (a, e, i, o, u) Example: Input: year Output: e Input: trust Output: r Input: qualification Output: u Input: defendant Output: 0.59
@@
word_unscrambling Convert the input to a word that is a common English word 0.62
Table 10: The best instruction discovered by ACING for all the 30 instruction-induction tasks using with Vicuna-13B as the white-box LLM.

### v2

@@ F.4 ACING with Budget Splitting
Spelling Letters_list 0.59 (0.02) 0.97 (0.03) 1.00 (0.00) 0.99 (0.01) 1.00 (0.00) 1.00 (0.00)
First_word_letter 0.00 (0.00) 1.00 (0.00) 1.00 (0.00) 1.00 (0.00) 1.00 (0.00) 1.00 (0.00)
Second_word_letter 0.00 (0.00) 0.63 (0.17) 0.35 (0.09) 0.39 (0.28) 0.40 (0.17) 0.70 (0.15)
Morpho-Syntax Singular_to_plural 1.00 (0.00) 1.00 (0.00) 0.99 (0.01) 0.95 (0.03) 0.99 (0.01) 0.95 (0.03)
Active_to_passive 1.00 (0.00) 0.99 (0.00) 0.98 (0.01) 1.00 (0.00) 1.00 (0.00) 1.00 (0.00)
Negation 0.79 (0.00) 0.84 (0.02) 0.65 (0.10) 0.58 (0.22) 0.82 (0.00) 0.71 (0.06)
Lexical Semantics Antonyms 0.79 (0.02) 0.70 (0.01) 0.76 (0.00) 0.84 (0.01) 0.76 (0.06) 0.74 (0.01)
Synonyms 0.14 (0.01) 0.19 (0.07) 0.22 (0.11) 0.19 (0.08) 0.12 (0.02) 0.13 (0.02)
Word_unscrambling 0.54 (0.00) 0.44 (0.06) 0.59 (0.06) 0.54 (0.02) 0.59 (0.05) 0.50 (0.07)
Phonetics Rhymes 0.59 (0.01) 0.52 (0.05) 0.99 (0.01) 0.36 (0.04) 0.60 (0.38) 0.57 (0.31)
Numerical Sum 0.87 (0.01) 1.00 (0.00) 1.00 (0.00) 0.70 (0.21) 0.98 (0.01) 1.00 (0.00)
Diff 0.00 (0.00) 0.99 (0.01) 1.00 (0.00) 0.93 (0.09) 0.97 (0.04) 1.00 (0.00)
Knowledge Larger_animal 0.72 (0.02) 0.58 (0.06) 0.63 (0.07) 0.81 (0.09) 0.86 (0.06) 0.84 (0.07)
Periodic_elements 0.99 (0.01) 0.92 (0.00) 0.96 (0.03) 1.00 (0.00) 0.98 (0.03) 0.98 (0.00)
Cognitive Tasks Cause_and_effect 0.44 (0.09) 0.48 (0.10) 0.52 (0.09) 0.55 (0.11) 0.76 (0.18) 0.69 (0.15)
Common_concept 0.03 (0.02) 0.17 (0.00) 0.14 (0.04) 0.09 (0.04) 0.10 (0.01) 0.19 (0.05)
Object_counting 0.30 (0.02) 0.50 (0.06) 0.38 (0.06) 0.40 (0.12) 0.48 (0.11) 0.41 (0.03)
Odd_one_out 0.32 (0.02) 0.64 (0.04) 0.57 (0.02) 0.25 (0.18) 0.59 (0.05) 0.64 (0.00)
Orthography_starts_with 0.23 (0.01) 0.47 (0.02) 0.41 (0.09) 0.54 (0.06) 0.54 (0.15) 0.60 (0.12)
Taxonomy_animal 0.02 (0.02) 0.38 (0.06) 0.67 (0.14) 0.85 (0.06) 0.53 (0.34) 0.71 (0.02)
Auto_categorization 0.31 (0.01) 0.20 (0.03) 0.29 (0.02) 0.07 (0.07) 0.27 (0.06) 0.29 (0.04)
Word_sorting 0.58 (0.01) 0.01 (0.00) 0.64 (0.05) 0.23 (0.20) 0.72 (0.02) 0.70 (0.03)
CLUE Sentence_similarity 0.00 (0.00) 0.05 (0.00) 0.10 (0.00) 0.00 (0.00) 0.13 (0.08) 0.13 (0.07)
Sentiment 0.90 (0.00) 0.63 (0.17) 0.88 (0.03) 0.88 (0.02) 0.88 (0.03) 0.89 (0.01)
Translation Num_to_verbal 0.13 (0.02) 1.00 (0.00) 0.99 (0.01) 1.00 (0.00) 1.00 (0.00) 0.99 (0.01)
Translation_en-de 0.83 (0.01) 0.80 (0.02) 0.82 (0.01) 0.77 (0.02) 0.82 (0.01) 0.82 (0.01)
Translation_en-es 0.86 (0.01) 0.76 (0.00) 0.67 (0.24) 0.89 (0.00) 0.86 (0.02) 0.87 (0.02)
Translation_en-fr 0.88 (0.01) 0.86 (0.00) 0.77 (0.06) 0.85 (0.02) 0.85 (0.02) 0.83 (0.01)
Style Informal_to_formal 0.57 (0.01) 0.50 (0.02) 0.48 (0.02) 0.54 (0.09) 0.44 (0.05) 0.44 (0.05)
Coding Auto_debugging 0.25 (0.00) 0.25 (0.00) 0.25 (0.00) 0.07 (0.07) 0.29 (0.07) 0.25 (0.00)
median score 0.49 0.63 0.66 0.64 0.76 0.71
# best-performing tasks 7 7 7 8 13 13
Table 15: Average test performance (and standard deviations) across 3 random seeds comparing ACING versus APE, InstructZero, EvoPrompt, and INSTINCT. The bottom rows report the median score and total number of best-performing tasks for each method.
F.5 ACING with Different Intrinsic (Action) Dimensions
@@
# best-performing tasks 8 10 10 15 14
@@
F.6 ACING with Different Number of Exemplars
@@
# best-performing tasks 11 24
@@
F.7 ACING with Different White-box models
@@
# best-performing tasks 15 21
@@ Appendix G Demonstrations with Human Instructions
To contextualize the performance of ACING, we compare its best-learned instructions against human-written instructions curated by Honovich et al. (2023). Table 19 presents a representative subset of tasks, categorized by linguistic and semantic attributes, along with input–output demonstrations, human instructions, and corresponding performance scores. While human instructions often perform strongly, ACING matches or exceeds them in the majority of cases, particularly on tasks like Antonyms, Rhymes, and Sentence Similarity, where learned instructions yield notable improvements. The comparison underscores ACING’s capacity not only to automate instruction crafting but also to outperform carefully designed human-written prompts across diverse task types. Summary statistics at the bottom of the table show that ACING achieves a higher average and median score, and wins on a greater number of tasks overall.
@@
Spelling First_word_letter cat $\rightarrow$ c Extract the first letter of the input word. 1.00 (0.00) 1.00 (0.00)
Second_word_letter cat $\rightarrow$ a Extract the second letter of the input word. 0.96 (0.00) 0.92 (0.00)
Letters_list cat $\rightarrow$ c a t Break the input word into letters, separated by spaces. 1.00 (0.00) 1.00 (0.00)
Morpho- syntax Singular_to_plural cat $\rightarrow$ cats Convert the input word to its plural form. 1.00 (0.00) 1.00 (0.00)
Active_to_passive The artist introduced the scientist. $\rightarrow$ The scientist was introduced by the artist. Write the input sentence in passive form. 1.00 (0.00) 1.00 (0.00)
Syntax Negation Time is finite $\rightarrow$ Time is not finite. Negate the input sentence. 0.81 (0.00) 0.82 (0.00)
Lexical Semantics Antonyms won $\rightarrow$ lost Write a word that means the opposite of the input word. 0.70 (0.00) 0.83 (0.00)
Synonyms alleged $\rightarrow$ supposed Write a word with a similar meaning to the input word. 0.14 (0.01) 0.13 (0.00)
Phonetics Rhymes sing $\rightarrow$ ring Write a word that rhymes with the input word. 0.61 (0.01) 1.00 (0.00)
Knowledge Larger_animal koala, snail $\rightarrow$ koala Write the larger of the two given animals. 0.94 (0.00) 0.94 (0.00)
Semantics Cause_and_effect Sentence 1: The soda went flat. Sentence 2: The bottle was left open. $\rightarrow$ The bottle was left open. Find which of the two given cause and effect sentences is the cause. 0.97 (0.02) 0.90 (0.02)
Common_concept guitars, pendulums, neutrinos $\rightarrow$ involve oscillations. Find a common characteristic for the given objects. 0.11 (0.01) 0.11 (0.00)
Style Informal_to_formal Please call once you get there $\rightarrow$ Please call upon your arrival. Rephrase the sentence in formal language. 0.63 (0.00) 0.50 (0.00)
Numerical Sum 22 10 $\rightarrow$ 32 Sum the two given numbers. 1.00 (0.00) 1.00 (0.00)
Diff 32 22 $\rightarrow$ 10 Subtract the second number from the first. 1.00 (0.00) 1.00 (0.00)
Num_to_Verbal 26 $\rightarrow$ twenty-six Write the number in English words. 1.00 (0.00) 1.00 (0.00)
Multi- lingual Translation_en-de game $\rightarrow$ Spiel Translate the word into German. 0.81 (0.00) 0.84 (0.00)
Translation_en-es game $\rightarrow$ juego Translate the word into Spanish. 0.89 (0.00) 0.88 (0.00)
Translation_en-fr game $\rightarrow$ jeu Translate the word into French. 0.86 (0.00) 0.87 (0.00)
GLUE Sentiment The film is small in scope, yet perfectly formed. $\rightarrow$ positive Determine whether a movie review is positive or negative. 0.89 (0.01) 0.91 (0.00)
Sentence_similarity Sentence 1: A man is smoking. Sentence 2: A man is skating. $\rightarrow$ 0 - definitely not Rate the semantic similarity of two input sentences on a scale of 0 - definitely not to 5 - perfectly. 0.00 (0.00) 0.21 (0.00)
median score 0.89 0.91
average score 0.78 0.80
@@
Table 19: Classified tasks into categories from the instruction-induction datasets. For each task, we provide a corresponding demonstration, with $\rightarrow$ separating the input from the output, along with its respective human instruction as proposed in Honovich et al. (2023). We tested these instructions, report their test scores (mean over 3 runs, standard deviation in parentheses), and compare them to our best test scores using ACING with Vicuna-13B as the white-box model.
Appendix H Assessing Instruction Clarity: Human and Readability Analyses
To assess the interpretability and accessibility of the instructions generated by ACING, we conduct both a human evaluation and an automated readability analysis.
H.1 Human Evaluation
We conducted a human annotation study across all instruction-induction tasks reported in Table˜2.
A total of 26 participants (none of whom are paper co-authors) volunteered to independently rate the clarity, coherence, and task faithfulness of instructions produced by ACING using a 5-point Likert scale. The following question was used to guide ratings:
Does the instruction clearly describe what is happening in the demonstration? Could a language model complete the task correctly by following this instruction alone?
Each instruction was presented alongside:
  • The task name and category
  • An input–output demonstration pair
  • The corresponding ACING-generated instruction
Results.
The results of the human evaluation are summarized below:
  • Maximum score: 4.92
  • Third quartile (Q3): 4.50
  • Median (Q2): 3.90
  • First quartile (Q1): 2.83
  • Mean score: 3.66  (SD = 1.02)
These results indicate that the majority of instructions are perceived as clear and well-aligned with the task by human annotators.
H.2 Automated Readability Analysis
To complement the human study, we applied three established readability formulas to quantify the linguistic accessibility of the 30 generated instructions in Table˜21:
  • Flesch Reading Ease (FRE) Flesch (1948): Scores range from 0 to 100, with higher scores indicating greater ease of reading.
  • Flesch-Kincaid Grade Level (FKG) Kincaid et al. (1975): Maps readability to U.S. school grade levels, with lower scores indicating simpler text.
  • Coleman-Liau Index (CLI) Coleman and Liau (1975): A character-based grade-level readability metric.
Findings.
Readability scores across instructions show the following trends:
  • Most instructions fall within the FRE range of 60–95, with a median 70.8, indicating that they are accessible to a general audience.
  • FKG scores mostly range between 3 and 9, with a median 7.0, consistent with middle school to early high school reading levels.
  • CLI scores mostly range between 2–10, with a median of 7.3. This aligns with the FKG analysis, indicating that instructions are generally suitable for readers with middle school to early high school reading levels.
Overall, these results confirm that ACING generates instructions that are not only effective—according to human judgment—but also accessible, as measured by standardized readability metrics. The full set of readability scores is included in Table˜20.
Instruction FRE $\uparrow$ FKG $\downarrow$ CLI $\downarrow$
Change the input to match the output, but the output is already in the passive voice 74.3 6.9 7.00
Take a word and change it to its opposite 94.3 2.3 2.18
Match the input to the output, and the answer is:\$Input\$: Nature Nanotechnology, Annual Review of Biochemistry, and The Lancet Neurology \$Output\$: top journals \$Input\$: Jeans, Tops, and Suits \$Output\$: Apparel 15.6 18.5 14.53
Input the code into a Python interpreter and observe the output 49.5 9.1 9.45
Find the sentence that is the cause of the effect in the pair of sentences 84.5 5.2 5.43
Make a connection between the input and output, but the connection is not clear 65.7 7.6 9.01
Find the difference between the two numbers 66.8 5.7 10.63
Create a function that takes a string as input and returns the first letter of the first word in the string 80.8 7.2 6.82
Convert the input into output using the same word order and with the same meaning 67.5 7.6 8.13
Create a program that takes two animals as input and outputs the animal that is bigger 58.4 9.1 8.09
Input the word “year” and the output was “y e a r” 96.0 2.9 -1.35
Flip the truth value of the statements in the input 86.7 3.7 5.60
Convert numbers to words 75.9 3.7 7.25
Provide a number that represents how many items are in the input 60.7 7.8 7.35
Find the word that does not belong in each group based on the given words 95.7 3.6 5.04
Find a word in the text that starts with the letter provided and to output that word 85.1 5.6 5.66
Find the name of the element based on its atomic number 72.6 5.9 5.24
Input the word that the program thought I was inputting and then output the word that program thought I was inputting 76.7 7.8 9.58
Input a word and output the letter that corresponds to the second letter in that word 69.0 7.6 7.72
Find a sentence pair that is probably not similar, and the output is 3 - probably 61.9 8.4 6.97
Classify each input as positive or negative based on the assessment of the corresponding movie 33.7 12.3 13.16
Add the suffix -s to the end of the word to make it plural 95.9 3.4 0.31
Find the sum of the two numbers 103.0 0.6 0.69
Input a word that is a synonym for the word that was output 83.0 4.9 2.89
Make the AI generate a sequence of animals based on the input provided 57.0 8.5 7.80
Provide a translation for each word in the English text into German 67.8 6.8 8.80
Translate the words from English to Spanish, but I noticed that some of the translations are not accurate 66.4 8.5 10.59
Create a program that would take an English word as input and output its French equivalent 63.7 8.4 9.54
Output the words in alphabetical order, but the output is not in alphabetical order 35.5 11.8 10.67
Convert the input to a word that is a common English word 81.9 4.8 3.97
Median 70.8 7.0 7.3
Table 20: Readability analysis of the best instructions generated by ACING, measured using Flesch Reading Ease (FRE) Flesch (1948), Flesch-Kincaid Grade Level (FKG) Kincaid et al. (1975), and Coleman-Liau Index (CLI) Coleman and Liau (1975). Higher FRE and lower FKG/CLI indicate easier-to-understand instructions.
Appendix I Our Best Learned Instructions
In this section, we present the best-learned instructions discovered by ACING for each of the 30 instruction induction tasks. These instructions were generated using Vicuna-13B as the white-box model and optimized to maximize performance on the corresponding black-box evaluation. Table 21 showcases the resulting instructions alongside their test scores, which reflect the accuracy or task-specific metric obtained on held-out examples. The diversity and clarity of the instructions demonstrate ACING’s ability to synthesize task-relevant, semantically grounded prompts that elicit strong responses from black-box LLMs. Notably, several tasks achieve perfect scores, while others expose task-specific challenges (e.g., common_concept and synonyms), highlighting the varying difficulty across the instruction spectrum.
@@ Appendix I Our Best Learned Instructions
second_word_letter Input a word and output the letter that corresponds to the second letter in that word 0.91
@@
word_unscrambling Convert the input to a word that is a common English word 0.63
Table 21: The best instruction discovered by ACING for all the 30 instruction-induction tasks using with Vicuna-13B as the white-box model.
Appendix J Use of AI Assistance.
We used AI assistants (e.g., ChatGPT) in a limited and supporting role during the preparation of this paper. Specifically, we used AI tools to assist with editing text for clarity and code debugging. All core ideas, algorithms, experiments, results, analyses, and technical writing were fully developed and executed by the authors. No AI system contributed to scientific decisions, modeling choices, or interpretation of findings.